OpenStudy (anonymous):

- 4.5 + 4.2 + 5.6 − 7.3 Explain this problem in steps.

OpenStudy (anonymous):

the answer is -2.

OpenStudy (anonymous):

thats wrong.

OpenStudy (studygurl14):

sorry i frogto the - in front of the first number. let me star over

OpenStudy (anonymous):

its okay.

OpenStudy (studygurl14):

-4.5 + 4.2 + 5.6 - 7.3 -4.5 + 4.2 = -0.3 -0.3 + 5.6 = 5.3 5.3 - 7.3 = -2

OpenStudy (anonymous):

correct!